pam is jean's older sister. the difference in their ages is 4. In 4 years the sum of their ages will be 36. Find the current ages

Found 2 solutions by ikleyn, MathTherapy:
Answer by ikleyn(52776)

.
In 4 years later we will have 


    Pam + Jean  = 36

     x  + (x-4) = 36.

    2x          = 36 + 4 = 40

     x                   = 40/2 = 20.


Hence, in 4 years later Pam will be 20 years old, while Jean will be 16 years old.


Hence, now  Palm is 16 years old and Jean is 12 years old.    ANSWER

Solved.



Answer by MathTherapy(10551)


pam is jean's older sister. the difference in their ages is 4. In 4 years the sum of their ages will be 36. Find the current ages